Use the intercept method to graph this equation: -4x + 2y = 8
a) What is the x-intercept(write as an ordered pair)?

b) What is the y-intercept(write as an ordered pair)?

c) Is the origin on the right or left side of the line?
1
Expert's answer
2020-02-27T11:41:03-0500

"-4x + 2y = 8"

a) What is the x-intercept

"\begin{cases}
 -4x+2y=8\\
 y=0
 \end{cases}" "=> -4x=8 =>x=-2\\=>x-intercept \, is \,A(-2,0)."

b) What is the y-intercept

"\begin{cases}
 -4x+2y=8\\
 x=0
 \end{cases}" "=>2y=8=>y=4=>\\y-intercept \, is \, B(0,4)."

c) Is the origin on the right or left side of the line?



take a look at graph, origin is on the right side of the line.